# Observation of Power Superbroadening of Spectral Line Profiles on IBM Quantum

### v2

@@ Abstract
While in high-resolution spectroscopy the push is for eliminating or even inverting the power broadening, there are applications where it is used to an advantage for it facilitates off-resonance excitation. Here, we present a number of shaped pulses that exhibit power broadening much greater than that of the rectangular pulse of the same pulse area. They are grouped in two families of pulse shapes. In particular, in regard to the width of the second Rabi oscillation maximum, the quadratic pulse family shows an increase by a factor of 3.3 whereas the even-exponent pulse family exhibits an increase by a factor of more than 3.5.
@@ I Introduction
Power broadening is a crucial phenomenon in quantum control, causing significant interaction between light and matter in a broader frequency spectrum away from resonance. While it is often associated with undesirable effects, such as reduced sensitivity in high-resolution spectroscopy and a cause for unwanted transitions in qubit control, there are applications where it is found to be beneficial. For instance, the same mechanism by which it activates unwanted transitions can be used to enable weak but desired transitions. One such application is laser cooling, where power broadening facilitates effective transitions even with large detunings  [1, 2, 3, 4]. Additionally, power broadening enables the preparation of certain quantum states in cases where necessary transitions are not accessible otherwise [5]. Finally, power broadening is widely used in spectroscopy to effectively increase the width of narrow resonance peaks that are difficult to detect, enabling their capture and accurate measurement [6, 7, 8, 9]. Power-broadened slowing of atoms and molecules plays a key role in robust loading of particles into magneto-optical traps (MOTs) [10, 11, 12, 13, 14]. By helping the interaction occur in a broader range of frequencies, the resonance requirement becomes less strict. Therefore, interactions that are otherwise hindered by the resonance condition are enabled.
@@
Off-resonant qubit excitations are affected by the temporal shape of the Rabi frequency envelope. Hence, the pulse shape properties have long been used to solve common problems in quantum systems, such as leakage, crosstalk, noise sensitivity etc [32, 31, 46, 47, 48, 39, 40, 41, 42, 43, 44, 45]. A promising technique for exerting control on the lineshape, pulse shaping allows to reduce or completely eliminate the power broadening of the qubit spectral line. Gaussian, Sine, Sine2 and other temporal pulse envelopes can be used to diminish it, while the sech pulse halts it completely [36, 24, 32]. Some earlier works even study the potential of pulse shape families that reverse the power broadening [37, 38, 31], introducing the concept of power narrowing using Lorentzian and Lorentzian-based temporally-shaped pulses.
@@ II Adiabatic Evolution in Two-State Systems
II.1 Two-State Schrödinger Equations
In a two-state quantum system, the time-dependent Schrödinger equation governs the evolution of the state vector $\mathbf{C}(t)\rangle$ given by the equation
@@ II.1 Two-State Schrödinger Equations
where $\mathbf{H}(t)$ is the time-dependent Hamiltonian of the system and $\mathbf{C}(t)=\left[c_{1}(t),c_{2}(t)\right]^{T}$ is the wavevector, consisting of $c_{1}(t)$ and $c_{2}(t)$ – the (complex) probability amplitudes of the two states.
@@ II.2 Hamiltonian in Two-State System
$\varepsilon_{\pm}(t)=\pm\frac{1}{2}\sqrt{\Delta^{2}(t)+\Omega^{2}(t)}.$ (3)
In the context of pulse shaping, both $\Delta$ and $\Omega$ can vary with time. Substituting this Hamiltonian into Eq. (1), we obtain the two-state Schrödinger system of equations
$2i\dot{c}_{1}(t)$ $=-\Delta(t)c_{1}(t)+\Omega(t)c_{2}(t),$
$2i\dot{c}_{2}(t)$ $=\Omega^{*}(t)c_{1}(t)+\Delta(t)c_{2}(t),$ (4)
@@ II.3 Adiabatic Evolution
Adiabatic evolution occurs when the system’s Hamiltonian changes slowly enough over time so the system remains in an instantaneous eigenstate of $H(t)$, provided it started in one. This implies that the system avoids transitions between eigenstates, evolving within the same eigenstate [50]. This abstinence from transitioning only occurs in the adiabatic basis, while in the bare basis, transitions between the ground and excited states may occur only in the presence of chirp, change in detuning, $\dot{\Delta}(t)$.
@@ II.4 Boosting power broadening – non-adiabaticity
The adiabatic condition requires the change in the Hamiltonian to be slow relative to the energy difference between the eigenstates. If this is satisfied, the system evolves adiabatically. In quantum power broadening and superbroadening, pulse shaping affects both the Rabi frequency $\Omega(t)$ and the adiabaticity conditions, making it possible to maintain control over transitions and achieve substantial broadening effects. This connection has been studied in relation to power narrowing in Ref. [31]. Expanding on the ideas in that study, we regard power broadening a consequence of the nonadiabatic nature of the pulse shape. Therefore, we demand significant non-adiabaticity to achieve strong power broadening. This is the reason we look for pulse shapes that have more non-adiabatic artifacts than the rectangular pulse and naturally violate the adiabatic condition
$\underbrace{\sqrt{\Omega^{2}(t)+\Delta^{2}(t)}}_{\varepsilon_{1}(t)}\gg\underbrace{\frac{\dot{\Omega}(t)\Delta(t)+\Omega(t)\dot{\Delta}(t)}{\Omega^{2}(t)+\Delta^{2}(t)}}_{\dot{\vartheta_{1}}(t)}.$ (5)
Developing on this notion, we notice that the larger $\dot{\Omega}(t)$ is, the less adiabatic the system is. Thus, we seek pulse shapes that exhibit sudden Rabi frequency changes concurrent with small Rabi frequencies.
The theory of adiabaticity begins with the construction of the adiabatic states — the (time-dependent) eigenstates of the Hamiltonian [44]. The original diabatic basis is related to the adiabatic one by the transformation matrix $\mathbf{R}(\vartheta_{1})$:
@@
$\mathbf{H}_{1}(t)$ $=\mathbf{R}(-\vartheta_{1})\mathbf{H}\mathbf{R}(\vartheta_{1})-i\hbar\mathbf{R}(-\vartheta_{1})\frac{d}{dt}\mathbf{R}(\vartheta_{1})$ (9)
@@
Figure 1: [Color online] The super-nonadiabatic coupling and superadiabatic splitting are shown for three different pulse shapes: (from left to right) a quadratic with $\beta=-1$, a quadratic with $\beta=1$ and an even-power pulse with $N=3$.
@@ II.5 Superadiabatic basis
The superadiabatic bases [44] can be used to compare the superadiabaticity of the current pulse shapes. The superadiabatic bases are obtained by repeated diagonalization of the previous-order superadiabatic bases. We regard the adiabatic basis an order-1 superadiabatic basis, and use it as a starting point in order to retrieve the $n=2$ basis. We begin with a phase transformation on the Hamiltonian,
@@
By repeating the process, we can also obtain the higher-order superadiabatic bases [44].
(a) Illustration of the face-changing quadratic pulse shapes with 6 different values of $\beta$: -1, 0, 0.25, 0.5, 0.75 and 1.
(b) Illustration of the observed excitation landscapes (transition probability against Rabi frequency amplitude and detuning) of face-changing quadratic pulse shapes with 6 different values of $\beta$: -1, 0, 0.25, 0.5, 0.75 and 1.
Figure 2: [Color online] Illustration of the excitation landscapes (bottom) with a legend of the used quadratic pulse shape (top).
(a) Illustration of the even exponent pulse shapes with 6 different values of $P$: 0, 1, 2, and 3.
(b) Illustration of the observed excitation landscapes (transition probability against Rabi freq. amplitude and detuning) of even exponent pulse shapes with 4 different values of $P$: 0, 1, 2, and 3.
Figure 3: [Color online] Illustration of the excitation landscapes (bottom) with a legend of the used even exponent pulse shape (top).
III The quadratic and the even exponent pulse shapes

### v3

@@ Abstract
While in high-resolution spectroscopy the push is for eliminating or even inverting the power broadening, there are applications where it is used to an advantage for it facilitates off-resonance excitation. Here, we present a number of shaped pulses that exhibit power broadening much greater than that of the rectangular pulse of the same pulse area. They are grouped in two families of pulse shapes. In particular, in regard to the width of the second Rabi oscillation maximum, the quadratic pulse family shows a spectral increase by a factor of 3.3 whereas the power-law pulse family exhibits an increase by a factor of more than 3.5.
@@ I Introduction
Power broadening is a well-known phenomenon in spectroscopy, causing significant interaction between light and matter in a broader frequency spectrum away from resonance. While it is often associated with undesirable effects, such as reduced sensitivity in high-resolution spectroscopy and a cause for unwanted transitions (cross talk) in qubit control, there are applications where it is found to be beneficial. For instance, the same mechanism by which it activates unwanted transitions can be used to enable weak but desired transitions. One such application is laser cooling, where power broadening facilitates effective transitions even with large detunings  [1, 2, 3, 4]. Additionally, power broadening enables the preparation of certain quantum states in cases where necessary transitions are not accessible otherwise [5]. Finally, power broadening is widely used in spectroscopy to effectively increase the width of narrow resonance peaks that are difficult to detect, enabling their capture and accurate measurement [6, 7, 8, 9]. Power-broadened slowing of atoms and molecules plays a key role in robust loading of particles into magneto-optical traps (MOTs) [10, 11, 12, 13, 14]. By helping the interaction occur in a broader range of frequencies, the resonance requirement becomes less strict. Therefore, interactions that are otherwise hindered by the resonance condition are enabled.
@@
Off-resonant qubit excitations are affected by the temporal shape of the Rabi frequency envelope. Hence, the pulse shape properties have long been used to solve common problems in quantum systems, such as leakage, crosstalk, noise sensitivity, etc. [32, 31, 47, 48, 49, 39, 40, 41, 42, 43, 44, 45, 46]. As a promising technique for exerting control on the lineshape, pulse shaping allows to reduce or completely eliminate the power broadening of the qubit spectral line. Power refers to the drive power at the qubit input. For fixed wiring and calibration, $P\propto\Omega_{0}^{2}$, so we report and control power via the Rabi frequency $\Omega_{0}/2\pi$. Gaussian, Sine, Sine2 and other temporal pulse envelopes can be used to diminish the broadening effect, while the sech pulse halts it completely [36, 24, 32]. Some earlier works even study the potential of pulse shape families that reverse the power broadening [37, 38, 31], introducing the concept of power narrowing using Lorentzian and Lorentzian-based temporally-shaped pulses.
@@ II Adiabatic Evolution in Two-State Systems
II.1 Two-State Schrödinger Equation
In a two-state quantum system, the time-dependent Schrödinger equation governs the evolution of the state vector $\mathbf{C}(t)$ by the equation (in natural units $\hbar=1$)
@@ II.1 Two-State Schrödinger Equation
where $\mathbf{H}(t)$ is the time-dependent Hamiltonian of the system and $\mathbf{C}(t)=\left[c_{1}(t),c_{2}(t)\right]^{T}$ is the state vector, consisting of $c_{1}(t)$ and $c_{2}(t)$ – the (complex) probability amplitudes of the two states.
@@ II.2 Hamiltonian in Two-State System
$\varepsilon_{\pm}(t)=\pm\frac{1}{2}\sqrt{\Omega^{2}(t)+\Delta^{2}(t)}.$ (3)
In the context of pulse shaping, both $\Delta(t)$ and $\Omega(t)$ can vary with time. Substituting this Hamiltonian into Eq. (1), we obtain the two-state Schrödinger system of equations
$i\dot{c}_{1}(t)$ $=-\frac{1}{2}\Delta(t)c_{1}(t)+\frac{1}{2}\Omega(t)c_{2}(t),$
$i\dot{c}_{2}(t)$ $=\frac{1}{2}\Omega^{*}(t)c_{1}(t)+\frac{1}{2}\Delta(t)c_{2}(t),$ (4)
@@ II.3 Adiabatic Evolution
Adiabatic evolution occurs when the system’s Hamiltonian changes slowly enough over time so the system remains in an instantaneous eigenstate of $\mathbf{H}(t)$, provided it started in one. This implies that the system avoids transitions between eigenstates, evolving within the same eigenstate [51]. This abstinence from transitioning only occurs in the adiabatic basis, while in the bare basis, transitions between the ground and excited states may occur only in the presence of chirp, change in detuning, $\dot{\Delta}(t)$.
@@ II.4 Boosting power broadening – non-adiabaticity
The adiabatic condition requires the change in the Hamiltonian to be slow relative to the energy difference between the eigenstates. If this is satisfied, the system evolves adiabatically. In quantum power broadening and power superbroadening — an enhanced form of power broadening which exceeds the broadening manifested by the rectangular pulse shape — pulse shaping affects both the Rabi frequency $\Omega(t)$ and the adiabaticity conditions, making it possible to maintain control over transitions and achieve substantial broadening effects. This connection has been studied in relation to power narrowing in Ref. [31]. Expanding on the ideas in that study, we regard power broadening as a consequence of the nonadiabatic nature of the pulse shape. Therefore, we demand significant non-adiabaticity to achieve strong power broadening. This is the reason we look for pulse shapes that have more non-adiabatic artifacts than the rectangular pulse and naturally violate the adiabatic condition
$\underbrace{\sqrt{\Omega^{2}(t)+\Delta^{2}(t)}}_{\varepsilon_{1}(t)}\gg\underbrace{\frac{\dot{\Omega}(t)\Delta(t)+\Omega(t)\dot{\Delta}(t)}{\Omega^{2}(t)+\Delta^{2}(t)}}_{\dot{\vartheta_{1}}(t)},$ (5)
where $\varepsilon_{1}(t)$ is the adiabatic splitting and $\dot{\vartheta}_{1}(t)$ is the nonadiabatic coupling. Whenever $\varepsilon_{1}(t)$ is larger than $\dot{\vartheta}_{1}(t)$ we have adiabatic evolution, described in Subsec. II.3. However, if $\varepsilon_{1}(t)$ is smaller than $\dot{\vartheta}_{1}(t)$, we see nonadiabatic effects such as highly off-resonant transitions and power broadening. Thus, we seek nonadiabatic pulse shapes that exhibit sudden Rabi frequency changes concurrent with small Rabi frequencies.
The theory of adiabaticity begins with the construction of the adiabatic states — the (time-dependent) eigenstates of the Hamiltonian [44, 51]. The original diabatic basis is related to the adiabatic one by the transformation matrix $\mathbf{R}(\vartheta_{1})$:
@@
$\mathbf{H}_{1}(t)$ $=\mathbf{R}(-\vartheta_{1})\mathbf{H}\mathbf{R}(\vartheta_{1})-i\mathbf{R}(-\vartheta_{1})\frac{d}{dt}\mathbf{R}(\vartheta_{1})$ (9)
@@
Figure 1: [Color online] Super-nonadiabatic coupling (dot-dashed curves) and superadiabatic energy splitting (dashed curves) for three drive envelopes (solid curves), shown left-to-right: (i) quadratic pulse with $\beta=–1$ (see Eq. (27)), (ii) quadratic pulse with $\beta=+1$ (see Eq. (27)), and (iii) power-law pulse with $P=3$ (see Eq. (28)). All traces share the same time axis and peak Rabi amplitude, allowing a direct comparison of how the sign of $\beta$ and the power-law exponent $P$ modify the nonadiabatic region near the pulse center.
@@ II.5 Superadiabatic basis
The superadiabatic (SA) bases [44, 52, 53, 54] can be used as a diagnostic to identify and quantify nonadiabatic regions generated by our pulse shape families. The (super)adiabatic bases may facilitate easier simulation when the evolution is (super)adiabatic and the corresponding condition is satisfied. In contrast, should one need to simulate the qubit dynamics when the evolution is nonadiabatic, it would be instead advantageous to perform exact numerical solution of the two-state Schrödinger equation, rather than use (super)adiabatic evolution. The superadiabatic bases are obtained by repeated diagonalization of the previous-order superadiabatic bases. We regard the adiabatic basis an order-1 superadiabatic basis, and use it as a starting point in order to retrieve the $n=2$ basis. We begin with a phase transformation on the Hamiltonian,
@@
By repeating the process, we can also obtain the higher-order superadiabatic bases [44, 52, 53, 54].
The super-nonadiabatic coupling and the SA splitting are plotted for three pulse shapes in Fig. 1 to help find nonadiabatic spots that predict the resulting power superbroadening.
II.6 Relation between adiabatic and superadiabatic conditions
Using the usual mixing angle from Eq. (8), its time derivative, the nonadiabatic coupling from Eq. (5), reads
$\dot{\vartheta}_{1}(t)=\frac{\dot{\Omega}(t)\,\Delta(t)-\Omega(t)\,\dot{\Delta}(t)}{2\,[\Omega^{2}(t)+\Delta^{2}(t)]}.$ (17)
The adiabatic condition (5) is $\varepsilon_{1}\gg|\dot{\vartheta}_{1}|.$ After one SA transformation, the splitting and the coupling are shown in Eq. (16):
$\varepsilon_{2}(t)=\sqrt{\varepsilon_{1}^{2}(t)+4\,\dot{\vartheta}_{1}^{2}(t)},$ (18)
$\dot{\vartheta}_{2}(t)=\frac{2\,\ddot{\vartheta}_{1}(t)\,\varepsilon_{1}(t)+2\,\dot{\vartheta}_{1}(t)\,\dot{\varepsilon}_{1}(t)}{\varepsilon_{1}^{2}(t)+4\,\dot{\vartheta}_{1}^{2}(t)}.$ (19)
For pulses characterized by a timescale $\tau$, the smoothness implies $|\dot{\Omega}|,|\dot{\Delta}|=\mathcal{O}(\tau^{-1})$ and $|\ddot{\Omega}|,|\ddot{\Delta}|=\mathcal{O}(\tau^{-2})$. Hence
$|\dot{\vartheta}_{1}|=\mathcal{O}\!\big((\varepsilon_{1}\tau)^{-1}\big),\qquad|\dot{\vartheta}_{2}|=\mathcal{O}\!\big(\varepsilon_{1}^{-3}\tau^{-2}\big).$ (20)
Since $\varepsilon_{2}\geq\varepsilon_{1}$, if (5) holds with margin $\gamma=\varepsilon_{1}/|\dot{\vartheta}_{1}|\gg 1$, then
$\frac{\varepsilon_{2}}{|\dot{\vartheta}_{2}|}\;\gtrsim\;\mathcal{O}(\gamma^{2}),$ (21)
so the SA condition (16) is not stricter than (5); it is typically easier to satisfy. Conversely, violating (16) marks a strongly nonadiabatic (“super-nonadiabatic”) regime where the SA coupling is large.
These observations motivate our envelope design principle: maximize $|\dot{\Omega}(t)|$ where $|\Omega(t)|$ is small (mid-pulse pits or sharp curvature). This maximizes $|\dot{\vartheta}_{1}|$ and $|\dot{\vartheta}_{2}|$ near the pulse center, producing the observed power superbroadening.
This asymptotic implication “adiabatic $\Rightarrow$ superadiabatic” refers to the regime where (5) holds with a large margin; it does not assert a pointwise ordering when the adiabatic condition is only marginally satisfied or violated.
For $\dot{\Delta}=0$,
$\dot{\vartheta}_{1}=\frac{\Delta\,\dot{\Omega}}{2(\Omega^{2}+\Delta^{2})}=\frac{\Delta\,\dot{\Omega}}{2\varepsilon_{1}^{2}},\qquad\varepsilon_{1}=\sqrt{\Omega^{2}+\Delta^{2}}.$ (22)
Hence the obstacle to adiabatic evolution depends essentially on the slope $|\dot{\Omega}|$. By contrast, the second–order coupling reads
$\dot{\vartheta}_{2}=\frac{2\,\ddot{\vartheta}_{1}\,\varepsilon_{1}+2\,\dot{\vartheta}_{1}\,\dot{\varepsilon}_{1}}{\varepsilon_{1}^{2}+4\dot{\vartheta}_{1}^{\,2}}=\frac{\Delta\big(\varepsilon_{1}^{2}\,\ddot{\Omega}-\Omega\,\dot{\Omega}^{\,2}\big)}{\varepsilon_{1}^{3}\big(\varepsilon_{1}^{2}+4\dot{\vartheta}_{1}^{\,2}\big)},$ (23)
where $\dot{\varepsilon}_{1}=\Omega\,\dot{\Omega}/\varepsilon_{1}$ and $\ddot{\vartheta}_{1}=\Delta(\varepsilon_{1}^{2}\,\ddot{\Omega}-2\,\Omega\,\dot{\Omega}^{\,2})/(2\,\varepsilon_{1}^{4})$. Near the pulse center of the chirp-free quadratic family, where $|\Omega|\ll|\Delta|$ so that $\varepsilon_{1}\simeq|\Delta|$, one obtains the form
$\dot{\vartheta}_{2}\;\simeq\;\mathrm{sgn}(\Delta)\;\frac{\,\ddot{\Omega}\;-\;(\Omega/|\Delta|^{2})\,\dot{\Omega}^{\,2}\,}{\,|\Delta|^{2}\Big(1+\frac{\dot{\Omega}^{\,2}}{|\Delta|^{4}}\Big)}\,.$ (24)
Thus the $4\dot{\vartheta}_{1}^{\,2}$ contribution is retained through the positive regularizing factor $1/(1+\dot{\Omega}^{\,2}/|\Delta|^{4})$. Using the adiabatic margin $\gamma_{1}=\varepsilon_{1}/|\dot{\vartheta}_{1}|\simeq 2|\Delta|^{2}/|\dot{\Omega}|$, we have
$\frac{\dot{\Omega}^{\,2}}{|\Delta|^{4}}=\frac{4}{\gamma_{1}^{2}},$ (25)
and therefore
$\dot{\vartheta}_{2}\;\simeq\;\mathrm{sgn}(\Delta)\;\frac{\,\ddot{\Omega}\;-\;(\Omega/|\Delta|^{2})\,\dot{\Omega}^{\,2}\,}{\,|\Delta|^{2}\,}\;\cdot\;\frac{\gamma_{1}^{2}}{\gamma_{1}^{2}+4}\,.$ (26)
Two regimes follow immediately:
(i) if $\gamma_{1}\gg 1$ and $|\Omega|\ll|\Delta|$, the denominator factor $\gamma_{1}^{2}/(\gamma_{1}^{2}+4)\to 1$ and $\dot{\vartheta}_{2}\simeq\mathrm{sgn}(\Delta)\,\ddot{\Omega}/|\Delta|^{2}$ (curvature–controlled);
(ii) if $\gamma_{1}=\mathcal{O}(1)$, the $4\dot{\vartheta}_{1}^{\,2}$ term cannot be neglected and it suppresses $\dot{\vartheta}_{2}$ by the factor $\gamma_{1}^{2}/(\gamma_{1}^{2}+4)$.
In both cases the denominator is positive, so the sign of $\dot{\vartheta}_{2}$ (and the $\beta=\pm 1$ contrast) is governed by the curvature term $\ddot{\Omega}$ instead of the slope $\dot{\Omega}$, while the denominator only rescales its magnitude.
Consequently, two chirp-free envelopes that mildly violate (or similarly satisfy) the adiabatic condition — because they have comparable $|\dot{\Omega}|$ — can still produce markedly different dynamics at second order due to opposite curvature near the midpoint. This explains the observed disparity between $\beta=-1$ and $\beta=+1$ and precisely shows the diagnostic value of the SA basis in our work. Furthermore, we can see in Ref. [44] how higher-order SA violations can be used to map the behavior of higher-exponent sinusoidal pulse shapes.
Figure 2: [Color online] Top: Quadratic pulse envelopes $\Omega(t)$ for $\beta\in\{-1,0,0.25,0.5,0.75,1\}$ at fixed duration. Bottom: Final excited-state population $P_{2}(T)$ versus detuning $\Delta/2\pi$ (x-axis) and Rabi amplitude $\Omega_{0}/2\pi$ (y-axis, $\propto\sqrt{\text{power}}$). The operational linewidth is defined as the full excitation width in $\Delta$ of the second Rabi maximum. Increasing $\beta$ enhances nonadiabatic coupling near the pulse midpoint (see Fig. 1), yielding pronounced power superbroadening relative to the rectangular case ($\beta=0$).
Figure 3: [Color online] Top: Power-law pulse envelopes $\Omega(t)$ with a mid-pulse pit of tunable width for $P=0,1,2,3$ (rectangular pulse at $P=0$). Bottom: Excitation landscapes. For $P>0$ the two horns act as Ramsey-like $\pi/2$ drives separated by a dark interval, yielding near-vertical interference fringes off-resonance. The pronounced increase of the operational linewidth is an indicator of power superbroadening for growing area.
III The quadratic and the power-law pulse shapes
We seek envelopes that produce large $|\dot{\vartheta}_{1}|$ (and, at second order, $|\dot{\vartheta}_{2}|$) at small $|\Omega|$, thereby enhancing nonadiabatic coupling away from resonance. The quadratic family flips curvature mid-pulse, isolating the role of slope sign ($\beta=-1$ vs. $\beta=+1$), while the power-law family enforces a mid-pulse pit of tunable width (parameter $P$), effectively creating two separated horns reminiscent of Ramsey $\pi/2$$\pi/2$ sequences. Both families
(i) keep $\Omega(t)\geq 0$ (no phase flips),
(ii) have a characteristic timescale $\tau$ given by $T/2$,
(iii) control non-adiabaticity with a single parameter,
(iv) include the rectangular pulse as a limit, and
(v) are easily applied under Qiskit amplitude/duration constraints, enabling a clean hardware demonstration. We then proceed to demonstrate the models’ excitation landscapes — post-pulse excited population versus detuning (x-axis) and Rabi frequency (y-axis) — on IBM’s ibm_sherbrooke (see specifics of the demonstration in Appendix A).